The volume of a rectangular solid is 810 meter cube. Find its dimensions if they are in the ratio 2:3:5.
Answer by lwsshak3(11628)

The volume of a rectangular solid is 810 meter cube. Find its dimensions if they are in the ratio 2:3:5.
**
let x=smallest of the three dimensions
1.5x=2nd dimension
2.5x=3rd dimension
..
x(1.5x)(2.5x)=810
3.75x^3=810
x^3=810/3.75=216
x=cube root of 216=6
1.5x=9
2.5x=15
check:
6*9*15=810